# Cosmic-ray tomography of shipping containers: A combination of complementary secondary particle and muon information using simulations

### v1

@@ I Introduction
This work focuses on the first direct comparison and statistical combination of the results acquired through MST and secondary particle analysis (SPA) building upon the methods presented in the previous work by the authors bib31 ; bib32 ; bib33 . At first, the setup of the simulation including the geometry of the detector and the container is briefly explained. Next, the separate reconstruction methodologies for SPA and MST are presented, which allow for the reconstruction and discrimination of the objects located inside the VOI. Following this, the combination procedure based on a clustering algorithm including the normalization and convolution of the reconstructed objects, as well as its segmentation over the noise-induced background is explained. In addition, the performance metrics used to evaluate the results before and after the combination are described. In conclusion, the results of this first-of-its-kind approach based on a consistent statistical combination of the complementary MST and SPA measurements are presented enabling further studies on such dual-channel analyses.
@@ II Simulation setup
The simulations are performed with the Geant4 Monte Carlo simulation toolkit bib34 ; bib35 to model all interactions between particles and matter. The precomputed and parametrized look-up tables of the Cosmic-ray Shower Library (CRY) bib12 are used to reproduce the cosmic-ray air showers and the kinematics of the resulting particles at sea level. For each simulated scene, 100 million air showers are generated, which is equivalent to a realistic scanning time of around 30 minutes.
@@
The target object consists of a 1 m3 cube, which is located in the middle of a simplified model of a shipping container made out of stainless steel. Five different scenes are generated, each with a different cube material: water, aluminum, vanadium, copper and lead. The origin of the coordinate system is set to be at the center of the container, with the x-axis along the width, the y-axis along the length and the z-axis along the height of the container. The detector configuration is based on a common design used in CRT and consists of four sets of three 50 mm thick Vinyltoluene based plastic scintillator layers with 10 cm spacing. As depicted in Figure 1, two of these sets are arranged vertically, one at the top and one at the bottom, while the other two are positioned on the left and right sides of the container. This configuration delineates a volume of interest with dimensions of 3.5 m × 7.0 m × 3.5 m.
@@
Figure 1: Visualization of the simulation setup. The detector layers are indicated by the blue layers surrounding the container. The reconstructed generic object positioned inside the container is represented by the colored dots. The trajectory of an air shower particle is represented by the purple line, while the pink lines show the paths of produced secondary particles.
@@
Particle information (e.g. energy, particle type, momentum direction) from each detector layer is directly retrieved from Geant4, as neither the particle identification, nor the readout from the scintillator material is emulated. Particles which don’t originate within the VOI or fail to intersect all three detection layers in any of the four detector sets are excluded from the analysis. The estimated trajectory of each particle is based on a least squares fitting method using the three detection positions across the three scintillator layers. For simplicity, perfect spatial resolution and a detection efficiency of 100 % are assumed as these values strongly correlate with the precise detector configuration, which is beyond the scope of this work. Further studies quantifying the impact of realistic detector conditions on the results of the SPA can be found in a previous work by the authors bib32 .
@@ III.1 Secondary particle analysis (SPA)
Initially, the VOI is discretized into a 3D voxel map. The reconstruction of the SPA measurements is based on the linear back-tracing of every detected photon, neutron and electron. This approach starts at the position at the innermost scintillator layer, back-traces the linear particle trajectory through the voxelized VOI and marks every voxel crossed. To assess the material characteristics per voxel, primarily the density and atomic number, the number of crossings per voxel is used as the representation value and referred to as the density score $s_{voxel}$. As the container and detector material also produce secondary particles and therefore act as a background source, the voxel map of an empty container reconstructed with 1 billion air shower events is subtracted for every scene.
In total, twelve voxel maps are measured for each scene, which correspond to specific detector and particle measurements defined in Table 1 and denoted $M1$-$M10$. Each measurement matches the combination of a given type of secondary particle (photon, neutron, electron), its energy and the different detectors (upper detector, sidewise detectors, lower detector). This allows a discrimination between secondary particles produced and air shower particles absorbed in the objects located inside the VOI. Further information on this scheme can be found in the previous work by the authors bib31 ; bib32 ; bib33 .
@@
The minimum density score threshold $t_{min}$ is chosen manually to enhance the reconstructed object over the background as much as possible while keeping the reconstructed shape and size of the object as close as possible to the original cube. The set of measurements and corresponding density score thresholds used to reconstruct the cube made out of water, aluminum, vanadium, copper and lead are given in Tables 2345 and 6 respectively. For all materials except lead, $M1$ and $M2.1$, as well as $M4$ and $M5.1$ are merged into one measurement to enhance the statistical significance of this voxel map during the combination procedure.
@@ III.2 Muon scattering tomography (MST)
Similar to the SPA, the VOI is also voxelized for the MST-based reconstruction. Two common MST algorithms are used in this work: PoCA bib36 and ASR bib37 . The Point-of-Closest-Approach (PoCA) method simplifies the multiple Coulomb scattering processes a muon encounters on its traversing of the examined volume by assuming that the scattering of the muon only happens at one point in the VOI. This so-called PoCA point is defined as the position in 3D space, where either the incoming and outgoing trajectory of the muon cross or where they are the closest to one another. The angle between the two trajectories is defined as the deflection of the muon and is assumed to reflect the material properties, mainly the density and atomic number. For all measured muons, this deflection angle gets assigned to the voxel corresponding to each PoCA point and for each voxel the median over all deflection angles assigned to this voxel gets referred to as the density score $s_{voxel}$.
@@ III.3 Muon and secondary particle combination (MST+SPA)
    Smoothing: A 3×3×3 convolution kernel is applied to the normalized and sharpened SPA and MST voxel maps. The kernel is configured such that the center voxel is assigned a weight of 100%, voxels sharing a face with the center receive a weight of 50%, those sharing an edge receive a weight of 35% and the corner voxels are weighted by 29%. These weights are dereived from the Euclidean distance from the kernel center. The smoothing operation ensures that statistical fluctuations in the results from SPA and MST are reduced facilitating a fair and consistent comparison and combination.
@@ IV Results
The results of the reconstruction of the water cube show overall a noticeable improvement for most of the metrics compared to the single MST and SPA results. The volumetric size of the reconstructed object is below the truth value, except for the ASR result. This is due to the significantly elongated length in the z-direction, which is a known artifact of the ASR algorithm and can be seen for all materials studied in this work. This also leads to a significantly higher Chamfer distance for ASR than for PoCA, SPA or the combined results. It can also be seen that the combined results are comparable for PoCA+SPA and ASR+SPA, which suggests that the artifacts in the ASR reconstruction can be effectively reduced by the combination with the SPA result.
For the scene with the aluminum cube, a similar noticeable improvement for most of the metrics can be seen as for the water cube scenario. However, the volumetric size of the reconstructed object is now matching the truth value for the combined measurements and SPA, but is higher for PoCA and ASR. While the average density score of the combined results is lower than for the water case, suggesting a less even distribution of the reconstructed density in the clustered object, $v_{reco}$ and $d_{c}$ show more similarity to the ground truth object.
@@
The reconstruction of the vanadium cube can also be improved with the combined approach in comparison to the single results. For this scene, the volumetric size of the reconstructed object is overall higher than the truth value for all measurements. From this and the other scenes it becomes evident that the SPA method results in the most even density score reconstruction, while PoCA is performing the poorest in this metric in general. This is also expected as the PoCA algorithm is known to show uneven density reconstruction due to its very simplified approach of assigning the scattering only to one voxel. PoCA and SPA show also clearly the highest similarity for the geometrical metrics compared to the ground truth object, while ASR is failing in these cases due to the already discussed artifact along the z-axis.
The combined results show an improvement in most of the metrics for the copper scene as well. The metrics are similar to the vanadium setup suggesting that the reconstruction methods are reaching their optimal working conditions for materials with medium densities. This can also be seen in the reconstructed shape along the z-direction for the ASR results, which is the worst for lower density material cubes.
Also for the highest density cube made out of lead, the combination results are showing an overall performance improvement for the majority of metrics compared to the single SPA, PoCA and ASR measurements. However, it is noticeable that with increasing density $s_{mean}$ drops continuously for the single, as well as for the combined results. This is likely the effect of various simplification assumptions, which are part of all reconstruction algorithms, leading to less continuous density distributions inside the reconstructed cube.
@@ V Conclusion
This work presented the first comparison and combination of stand-alone muon and secondary particle results in the field of cosmic-ray tomography by analyzing the contents of shipping container based on simulations. The results demonstrate that incorporating secondary particle information via this novel dual-channel analysis technique can serve as a valuable complementary data source. Noticeable improvements over single MST and SPA results are achieved using a principled algorithmic combination procedure, however utilizing a simplified detection and geometric assumptions. Yet, these improvements persist across a range of materials and densities, thus highlighting the applicability of the proposed method.

### v2

@@ I Introduction
This work focuses on the first direct comparison and statistical combination of the results acquired through MST and secondary particle analysis (SPA) building upon the methods presented in the previous work by the authors bib31 ; bib32 ; bib33 , which are the first of its kind for a stand-alone secondary analysis approach. At first, the setup of the simulation including the geometry of the detector and the container is briefly explained. Next, the separate reconstruction methodologies for SPA and MST are presented, which allow for the reconstruction and discrimination of the objects located inside the VOI. Following this, the combination procedure based on a clustering algorithm including the normalization and convolution of the reconstructed objects, as well as its segmentation over the noise-induced background is explained. In addition, the performance metrics used to evaluate the results before and after the combination are described. In conclusion, the results of this first-of-its-kind approach based on a consistent statistical combination of the complementary MST and SPA measurements are presented enabling further studies on such dual-channel analyses.
@@ II Simulation setup
The simulations are performed with the Geant4 Monte Carlo simulation toolkit bib34 ; bib35 to model all interactions between particles and matter. The precomputed and parametrized look-up tables of the Cosmic-ray Shower Library (CRY) bib12 are used to reproduce the cosmic-ray air showers and the kinematics of the resulting particles at sea level. The air showers are set to contain muons, photons, electrons, protons, neutrons and pions, as well as a maximum number of 30 particles within an air shower. In addition, the side length of the generation plane is 10 m and the latitude is set to 42 degrees north. For each simulated scene, 100 million air showers are generated, which is equivalent to a scanning time of around 30 minutes, which is realistic under regular operation conditions for a secondary custom control.
@@
The target object consists of a 1 m3 cube, which is located in the middle of a simplified model of a shipping container made out of stainless steel. The size and shape of the target object was chosen to imitate standard cargo objects inside a container, such as intermediate bulk containers or loaded europallets. Five different scenes are generated, each with a different cube material: water, aluminum, vanadium, copper and lead. The origin of the coordinate system is set to be at the center of the container, with the x-axis along the width, the y-axis along the length and the z-axis along the height of the container. The detector configuration is based on a common design used in CRT bib08 ; bib14 ; bib26 ; bib29 and consists of four sets of three 50 mm thick polyvinyl toluene based plastic scintillator layers with 10 cm spacing. As depicted in Figure 1, two of these sets are arranged vertically, one at the top and one at the bottom, while the other two are positioned on the left and right sides of the container. This configuration delineates a volume of interest with dimensions of 3.5 m × 7.0 m × 3.5 m.
@@
Figure 1: Visualization of the simulation setup. The detector layers are indicated by the blue planes surrounding the container, in which the target object is located. The trajectory of an air shower particle is represented by the green line, while the pink lines show the paths of secondary particles produced during the depicted interaction of the air shower particle with the target cube.
Particle information (e.g. energy, particle type, momentum direction) from each detector layer is directly retrieved from Geant4, as neither the particle identification, nor the readout from the scintillator material is emulated. The goal of this work is the development and evaluation of the theoretical groundwork for a first stand-alone secondary particle analysis. Therefore, all studies are conducted on a generalized level without a dedicated detector design in place. The main reason of the placement of the plastic scintillator material is the emulation of the impact of a sensitive detection material on the secondary particle kinematics.
@@
Particles which don’t originate within the VOI or fail to intersect all three detection layers in any of the four detector sets are excluded from the analysis. The estimated trajectory of each particle (muon, photon, neutron and electron) is based on a least squares fitting method using the three detection positions across the three scintillator layers. For simplicity, perfect spatial resolution and a detection efficiency of 100 % are assumed as these values strongly correlate with the precise detector configuration, which is beyond the scope of this work. Further studies quantifying the impact of realistic detector conditions on the results of the SPA can be found in a previous work by the authors bib32 .
@@ III.1 Secondary particle analysis (SPA)
Initially, the VOI is discretized into a 3D voxel map. The reconstruction of the SPA measurements is based on the linear back-tracing of every detected photon, neutron and electron. This assumption can be justified by the fact that most secondary particles reaching the detector are produced in the peripheral areas of the cube inside the container. Particles originating from the inner areas are less likely to leave the volume of the cube due to self-shielding effects. Since no other stand-alone reconstruction approach for secondary particles exists, the presented method is the first of its kind. The proposed approach starts at the position at the innermost scintillator layer, back-traces the linear particle trajectory through the voxelized VOI and marks every voxel crossed. To assess the material characteristics per voxel, primarily the density and atomic number, the number of crossings per voxel is used as the representation value and referred to as the density score $s_{voxel}$. As the container and detector material also produce secondary particles and therefore act as a background source, the voxel map of an empty container reconstructed with 1 billion air shower events is subtracted for every scene.
In total, twelve voxel maps are measured for each scene, which correspond to specific detector and particle measurements defined in Table 1 and denoted $M1$-$M10$. Each measurement matches the combination of a given type of secondary particle (photon, neutron, electron), its energy and the different detectors (upper detector, sidewise detectors, lower detector). This allows a discrimination between secondary particles produced and air shower particles, as well as secondary particles absorbed in the objects located inside the VOI, which is mainly achieved by separating particles with low energy from particles with high energy.
Each measurements contains a unique set of information about the object properties. For instance, low-density objects emit and absorb a significantly lower amount of neutrons than high-density materials, while the detection of low-energy secondary photons is related to the presence of a low-density material due to the increased probability of self-shielding for high-density objects. Furthermore, the absorption of air shower particles can only be measured with the lower detector, while the production of secondary particles is mainly prominent in the upper and sidewise detectors. Further information on this scheme can be found in the previous work by the authors bib31 ; bib32 ; bib33 .
@@
The minimum density score threshold $t_{min}$ is chosen manually for each measurement to enhance the reconstructed object over the background as much as possible while keeping the reconstructed shape and size of the object as close as possible to the original cube. The application of such a threshold is necessary as the back-tracing approach leads to a large level of noise around the reconstructed object in each voxel map. In order to remove this noise related background, a voxel is only considered for the final combined voxel map if it fulfills the threshold requirement in all of the measurements selected for the combination. The dependency of the threshold on the maximum density score was chosen to allow for a material dependent tuning of this parameter in order to improve the material discrimination.
The set of measurements and corresponding density score thresholds used to reconstruct the cube made out of water, aluminum, vanadium, copper and lead are given in Tables 2345 and 6 respectively. The density score thresholds slowly increase for all measurements with increasing density as the reconstructed cube becomes more significant over the background due to a higher production rate of secondary particles. Neutrons in the upper and sidewise detectors are only relevant for the lead cube due to the low probability of induced fission for low and medium density materials. For all materials except lead, $M1$ and $M2.1$, as well as $M4$ and $M5.1$ are merged into one measurement to enhance the statistical significance of this voxel map during the combination procedure.
@@ III.2 Muon scattering tomography (MST)
Similar to the SPA, the VOI is also voxelized for the MST-based reconstruction. Two common MST algorithms are used in this work: PoCA bib36 and ASR bib37 . The Point-of-Closest-Approach (PoCA) method simplifies the multiple Coulomb scattering processes a muon encounters on its traversing of the examined volume by assuming that the scattering of the muon only happens at one point in the VOI. This so-called PoCA point is defined as the position in 3D space, where either the incoming and outgoing trajectory of the muon cross or where they are the closest to one another. The angle between the two trajectories is defined as the deflection of the muon and is assumed to reflect the material properties, mainly the density and atomic number bib13 ; bib14 ; bib21 . For all measured muons, this deflection angle gets assigned to the voxel corresponding to each PoCA point and for each voxel the median over all deflection angles assigned to this voxel gets referred to as the density score $s_{voxel}$.
@@ III.3 Muon and secondary particle combination (MST+SPA)
    Smoothing: A 3×3×3 convolution kernel is applied to the normalized and sharpened SPA and MST voxel maps. The kernel is configured such that the center voxel is assigned a weight of 100%, voxels sharing a face with the center receive a weight of 50%, those sharing an edge receive a weight of 35% and the corner voxels are weighted by 29%. These weights are derived from the Euclidean distance from the kernel center. The smoothing operation ensures that statistical fluctuations in the results from SPA and MST are reduced facilitating a fair and consistent comparison and combination.
@@ IV Results
The results of the reconstruction of the water cube in Table 8 show overall a noticeable improvement for most of the metrics compared to the single MST and SPA results. The volumetric size of the reconstructed object is below the ground-truth value of 1.21 m3, except for the ASR result, which overestimates the volume at 1.89 m3. This is due to the significantly elongated length in the z-direction, which is a known artifact of the ASR algorithm and can be seen for all materials studied in this work. This also leads to a significantly higher Chamfer distance of 44.3 for ASR compared to the other algorithms. It can also be seen that the combined results are comparable for PoCA+SPA and ASR+SPA, which suggests that the artifacts in the ASR reconstruction can be effectively reduced by the combination with the SPA result.
For the scene with the aluminum cube, a similar noticeable improvement for most of the metrics in Table 9 can be seen similar to the water cube scenario. However, the volumetric size of the reconstructed object is now matching the truth value for the combined measurements and SPA, but is higher for PoCA and ASR. While the average density score of the combined results is lower than for the water case, suggesting a less even distribution of the reconstructed density in the clustered object, $v_{reco}$ and $d_{c}$ show more similarity to the ground truth object.
@@
The reconstruction of the vanadium cube can also be improved with the combined approach in comparison to the single results as shown in Table 10. For this scene, the volumetric size of the reconstructed object is overall higher than the truth value for all measurements. In addition, it becomes evident that compared to PoCA and ASR the SPA method shows the best performance in reconstructing a high density score, which is related to an even density score distribution of the cube, while PoCA is performing the poorest in this metric in general. This is also expected as the PoCA algorithm is known to show uneven density reconstruction due to its very simplified approach of assigning the scattering only to one voxel. PoCA and SPA show also clearly the highest similarity for the geometrical metrics compared to the ground truth object, while ASR is failing in these cases due to the already discussed artifact along the z-axis.
The combined results in Table 11 show an improvement in most of the metrics even for the copper scene. The metrics are similar to the vanadium setup suggesting that the reconstruction methods are reaching their optimal working conditions for materials with medium densities. This can also be seen in the reconstructed shape along the z-direction for the ASR results, which is the worst for lower density material cubes.
Also for the highest density cube made out of lead, the combination results are showing an overall performance improvement for the majority of metrics compared to the single SPA, PoCA and ASR measurements as shown in Table 12. However, it is noticeable that with increasing density $s_{mean}$ drops continuously for the single, as well as for the combined results. This is likely the effect of various simplification assumptions, which are part of all reconstruction algorithms, leading to less continuous density distributions inside the reconstructed cube.
@@ V Conclusion
This work presented the first comparison and combination of stand-alone muon and secondary particle results in the field of cosmic-ray tomography by analyzing the contents of a shipping container with a simplified detection and geometry setup. The simulation-based study evaluated the performance of the PoCA, ASR and SPA reconstruction methods, as well as the combinations of PoCA+SPA and ASR+SPA to reconstruct a 1 m3 cube made out of five different materials inside a shipping container. In order to ensure a consistent comparison and combination between the different methods, the study follows a multi-step procedure to segment the reconstructed cube over the background. The assessment of the performance of the different reconstruction algorithms is based on dedicated properties of the reconstructed object, such as its density score and other geometric quantities.
The results demonstrate that incorporating secondary particle information via this novel dual-channel analysis technique can serve as a valuable complementary data source. Noticeable improvements for all performance metrics over single MST and SPA results are achieved reducing known reconstruction artifacts of the utilized reconstruction algorithms. The improvements persist across a range of materials and densities, thus highlighting the applicability of the proposed method.